# Conceptualization of this study, Methodology, Software, Writing - original draft

### v1

@@ 5.1.2 Evaluation Metrics
Dataset Metric GRU4Rec NARM STAMP Caser SASRec BERT4Rec
w/o w w/o w w/o w w/o w w/o w w/o w
ML-100k HR@5 0.0191 0.0723 0.0180 0.0756 0.0201 0.0546 0.0212 0.0594 0.0191 0.0682 0.0191 0.0708
HR@10 0.0286 0.1249 0.0403 0.1317 0.0392 0.0943 0.0339 0.1051 0.0371 0.1203 0.0414 0.1270
HR@20 0.0594 0.2169 0.0657 0.2202 0.0700 0.1702 0.0679 0.1916 0.0764 0.2123 0.0912 0.2126
NDCG@5 0.0104 0.0453 0.0132 0.0470 0.0115 0.0331 0.0113 0.0362 0.0114 0.0437 0.0117 0.0449
NDCG@10 0.0134 0.0621 0.0202 0.0650 0.0176 0.0459 0.0153 0.0509 0.0172 0.0603 0.0189 0.0631
NDCG@20 0.0212 0.0821 0.0267 0.0871 0.0253 0.0648 0.0238 0.0726 0.0270 0.0834 0.0315 0.0845
MRR@20 0.0109 0.0494 0.0162 0.0509 0.0132 0.0365 0.0119 0.0409 0.0139 0.0487 0.0157 0.0498
Beauty HR@5 0.0077 0.0295 0.0120 0.0365 0.0080 0.0421 0.0072 0.0196 0.0242 0.0436 0.0060 0.0522
HR@10 0.0135 0.0481 0.0209 0.0565 0.0135 0.0638 0.0133 0.0340 0.0386 0.0632 0.0127 0.0714
HR@20 0.0256 0.0753 0.0367 0.0855 0.0231 0.0932 0.0235 0.0550 0.0561 0.0881 0.0204 0.0955
NDCG@5 0.0045 0.0194 0.0071 0.0243 0.0046 0.0285 0.0044 0.0120 0.0129 0.0307 0.0037 0.0378
NDCG@10 0.0064 0.0254 0.0099 0.0307 0.0064 0.0355 0.0064 0.0166 0.0175 0.0370 0.0059 0.0439
NDCG@20 0.0094 0.0322 0.0139 0.0380 0.0088 0.0429 0.0090 0.0219 0.0219 0.0434 0.0078 0.0500
MRR@20 0.0051 0.0204 0.0077 0.0249 0.0049 0.0289 0.0051 0.1281 0.0122 0.0307 0.0044 0.0372
Sports HR@5 0.0064 0.0176 0.0099 0.0199 0.0071 0.0260 0.0069 0.0119 0.0113 0.0258 0.0055 0.0271
HR@10 0.0114 0.0293 0.0138 0.0327 0.0123 0.0413 0.0115 0.0199 0.0175 0.0377 0.0104 0.0387
HR@20 0.0183 0.0465 0.0223 0.0517 0.0182 0.0629 0.0178 0.0329 0.0268 0.0546 0.0167 0.0553
NDCG@5 0.0035 0.0115 0.0058 0.0127 0.0046 0.0169 0.0046 0.0078 0.0059 0.0183 0.0036 0.0192
NDCG@10 0.0051 0.0153 0.0073 0.0169 0.0062 0.0218 0.0061 0.0103 0.0079 0.0221 0.0051 0.0229
NDCG@20 0.0068 0.0196 0.0094 0.0216 0.0077 0.0272 0.0077 0.0136 0.0102 0.0264 0.0067 0.0271
MRR@20 0.0036 0.0122 0.0059 0.0134 0.0048 0.0174 0.0049 0.0083 0.0055 0.0185 0.0040 0.0192
Yelp HR@5 0.0057 0.0215 0.0113 0.0248 0.0060 0.0234 0.0045 0.0224 0.0293 0.0212 0.0087 0.0243
HR@10 0.0102 0.0375 0.0187 0.0416 0.0099 0.0390 0.0084 0.0336 0.0352 0.0357 0.0159 0.0418
HR@20 0.0184 0.0631 0.0315 0.0686 0.0161 0.0645 0.0146 0.0517 0.0439 0.0584 0.0273 0.0695
NDCG@5 0.0034 0.0133 0.0075 0.0157 0.0038 0.0152 0.0028 0.0158 0.0251 0.0136 0.0054 0.0151
NDCG@10 0.0048 0.0184 0.0099 0.0211 0.0051 0.0202 0.0040 0.0194 0.0270 0.0183 0.0077 0.0207
NDCG@20 0.0068 0.0248 0.0131 0.0278 0.0066 0.0266 0.0055 0.0239 0.0292 0.0240 0.0105 0.0277
MRR@20 0.0037 0.0144 0.0081 0.0167 0.0040 0.0162 0.0031 0.0163 0.0250 0.0146 0.0060 0.0163
ML-1M HR@5 0.0194 0.1689 0.0151 0.1676 0.0232 0.1464 0.0104 0.1394 0.0397 0.1455 0.0224 0.1351
HR@10 0.0373 0.2473 0.0349 0.2457 0.0440 0.2145 0.0215 0.2158 0.0666 0.2251 0.0495 0.2095
HR@20 0.0690 0.3445 0.0591 0.3430 0.0677 0.3010 0.0589 0.3173 0.1007 0.3260 0.0980 0.3062
NDCG@5 0.0135 0.1139 0.0080 0.1124 0.0150 0.0993 0.0063 0.0899 0.0207 0.0936 0.0132 0.0873
NDCG@10 0.0190 0.1392 0.0144 0.1376 0.0218 0.1212 0.0099 0.1145 0.0294 0.1192 0.0218 0.1112
NDCG@20 0.0270 0.1636 0.0205 0.1621 0.0278 0.1430 0.0194 0.1401 0.0379 0.1447 0.0339 0.1356
MRR@20 0.0159 0.1129 0.0100 0.1113 0.0168 0.0988 0.0091 0.0908 0.0203 0.0940 0.0169 0.0881
@@ Acknowledgement
This work was supported by the National Natural Science Foundation of China [grant number 62141201]; the Natural Science Foundation of Chongqing, China [grant number CSTB2022NSCQ-MSX1672]; the Major Project of Science and Technology Research Program of Chongqing Education Commission of China [grant number KJZD-M202201102].

### v2

@@ 5.1.2 Evaluation Metrics
Dataset Metric GRU4Rec NARM STAMP Caser SASRec CL4SRec BERT4Rec
w/o w w/o w w/o w w/o w w/o w w/o w w/o w
ML-100k HR@5 0.0191 0.0723 0.0180 0.0756 0.0201 0.0546 0.0212 0.0594 0.0191 0.0682 0.0289 0.0437 0.0191 0.0708
HR@10 0.0286 0.1249 0.0403 0.1317 0.0392 0.0943 0.0339 0.1051 0.0371 0.1203 0.0579 0.0777 0.0414 0.1270
HR@20 0.0594 0.2169 0.0657 0.2202 0.0700 0.1702 0.0679 0.1916 0.0764 0.2123 0.1107 0.1404 0.0912 0.2126
NDCG@5 0.0104 0.0453 0.0132 0.0470 0.0115 0.0331 0.0113 0.0362 0.0114 0.0437 0.0174 0.0267 0.0117 0.0449
NDCG@10 0.0134 0.0621 0.0202 0.0650 0.0176 0.0459 0.0153 0.0509 0.0172 0.0603 0.0265 0.0378 0.0189 0.0631
NDCG@20 0.0212 0.0821 0.0267 0.0871 0.0253 0.0648 0.0238 0.0726 0.0270 0.0834 0.0398 0.0536 0.0315 0.0845
MRR@20 0.0109 0.0494 0.0162 0.0509 0.0132 0.0365 0.0119 0.0409 0.0139 0.0487 0.0209 0.0301 0.0157 0.0498
Beauty HR@5 0.0077 0.0295 0.0120 0.0365 0.0080 0.0421 0.0072 0.0196 0.0242 0.0436 0.0208 0.0249 0.0060 0.0522
HR@10 0.0135 0.0481 0.0209 0.0565 0.0135 0.0638 0.0133 0.0340 0.0386 0.0632 0.0379 0.0434 0.0127 0.0714
HR@20 0.0256 0.0753 0.0367 0.0855 0.0231 0.0932 0.0235 0.0550 0.0561 0.0881 0.0618 0.0694 0.0204 0.0955
NDCG@5 0.0045 0.0194 0.0071 0.0243 0.0046 0.0285 0.0044 0.0120 0.0129 0.0307 0.0116 0.0156 0.0037 0.0378
NDCG@10 0.0064 0.0254 0.0099 0.0307 0.0064 0.0355 0.0064 0.0166 0.0175 0.0370 0.0171 0.0215 0.0059 0.0439
NDCG@20 0.0094 0.0322 0.0139 0.0380 0.0088 0.0429 0.0090 0.0219 0.0219 0.0434 0.0231 0.0280 0.0078 0.0500
MRR@20 0.0051 0.0204 0.0077 0.0249 0.0049 0.0289 0.0051 0.0128 0.0122 0.0307 0.0125 0.0167 0.0044 0.0372
Sports HR@5 0.0064 0.0176 0.0099 0.0199 0.0071 0.0260 0.0069 0.0119 0.0113 0.0258 0.0150 0.0270 0.0055 0.0271
HR@10 0.0114 0.0293 0.0138 0.0327 0.0123 0.0413 0.0115 0.0199 0.0175 0.0377 0.0256 0.0422 0.0104 0.0387
HR@20 0.0183 0.0465 0.0223 0.0517 0.0182 0.0629 0.0178 0.0329 0.0268 0.0546 0.0418 0.0638 0.0167 0.0553
NDCG@5 0.0035 0.0115 0.0058 0.0127 0.0046 0.0169 0.0046 0.0078 0.0059 0.0183 0.0088 0.0177 0.0036 0.0192
NDCG@10 0.0051 0.0153 0.0073 0.0169 0.0062 0.0218 0.0061 0.0103 0.0079 0.0221 0.0122 0.0226 0.0051 0.0229
NDCG@20 0.0068 0.0196 0.0094 0.0216 0.0077 0.0272 0.0077 0.0136 0.0102 0.0264 0.0163 0.0280 0.0067 0.0271
MRR@20 0.0036 0.0122 0.0059 0.0134 0.0048 0.0174 0.0049 0.0083 0.0055 0.0185 0.0093 0.0182 0.0040 0.0192
Yelp HR@5 0.0057 0.0215 0.0113 0.0248 0.0060 0.0234 0.0045 0.0224 0.0293 0.0212 0.0276 0.0307 0.0087 0.0243
HR@10 0.0102 0.0375 0.0187 0.0416 0.0099 0.0390 0.0084 0.0336 0.0352 0.0357 0.0425 0.0471 0.0159 0.0418
HR@20 0.0184 0.0631 0.0315 0.0686 0.0161 0.0645 0.0146 0.0517 0.0439 0.0584 0.0657 0.0711 0.0273 0.0695
NDCG@5 0.0034 0.0133 0.0075 0.0157 0.0038 0.0152 0.0028 0.0158 0.0251 0.0136 0.0185 0.0209 0.0054 0.0151
NDCG@10 0.0048 0.0184 0.0099 0.0211 0.0051 0.0202 0.0040 0.0194 0.0270 0.0183 0.0233 0.0261 0.0077 0.0207
NDCG@20 0.0068 0.0248 0.0131 0.0278 0.0066 0.0266 0.0055 0.0239 0.0292 0.0240 0.0291 0.0321 0.0105 0.0277
MRR@20 0.0037 0.0144 0.0081 0.0167 0.0040 0.0162 0.0031 0.0163 0.0250 0.0146 0.0191 0.0214 0.0060 0.0163
ML-1M HR@5 0.0194 0.1689 0.0151 0.1676 0.0232 0.1464 0.0104 0.1394 0.0397 0.1455 0.0369 0.0754 0.0224 0.1351
HR@10 0.0373 0.2473 0.0349 0.2457 0.0440 0.2145 0.0215 0.2158 0.0666 0.2251 0.0684 0.1294 0.0495 0.2095
HR@20 0.0690 0.3445 0.0591 0.3430 0.0677 0.3010 0.0589 0.3173 0.1007 0.3260 0.1204 0.2064 0.0980 0.3062
NDCG@5 0.0135 0.1139 0.0080 0.1124 0.0150 0.0993 0.0063 0.0899 0.0207 0.0936 0.0216 0.0471 0.0132 0.0873
NDCG@10 0.0190 0.1392 0.0144 0.1376 0.0218 0.1212 0.0099 0.1145 0.0294 0.1192 0.0317 0.0644 0.0218 0.1112
NDCG@20 0.0270 0.1636 0.0205 0.1621 0.0278 0.1430 0.0194 0.1401 0.0379 0.1447 0.0447 0.0837 0.0339 0.1356
MRR@20 0.0159 0.1129 0.0100 0.1113 0.0168 0.0988 0.0091 0.0908 0.0203 0.0940 0.0243 0.0501 0.0169 0.0881
@@ 5.1.3 Competing Models
  • CL4SRec (Xie et al., 2022). It is a contrastive learning based sequential recommender which attempts to learn better sequence representations by leveraging the contrastive learning framework to obtain self-supervision signals.
@@ 5.6 Impact of Training Set Proportion
Figure 7: Performance Comparison of our model with the Gumbel-softmax strategy and the clipping strategy on the ML-100k dataset.
5.7 Impact of the Gumbel-softmax Strategy
To validate the effectiveness of incorporating the Gumbel-softmax strategy in our proposed method, we replace it with a simple clipping strategy. We vary the value of the clipping threshold from 0.1 to 0.9 with a step size of 0.1, and the performance on the ML-100k dataset is demonstrated in Figure 7. We can observe that the model performance with the clipping strategy improves if a higher threshold value is used and reaches a peak when the threshold value is 0.2. Then the performance starts to drop continuously as the threshold value increases. In addition, we can clearly observe that the model performance with the Gumbel-softmax strategy is consistently superior to that of the clipping strategy, which suggests that it is beneficial to employ the Gumbel-softmax strategy in our proposed method. This is probably because the clipping strategy always obtain the same output while the Gumbel-softmax strategy might get probability output due to the incorporation of the Gumbel noise.
@@ Acknowledgement
This work was supported by the National Natural Science Foundation of China [grant number 62141201]; the Natural Science Foundation of Chongqing, China [grant number CSTB2022NSCQ-MSX1672]; the Chongqing Talent Plan Project, China [grant number CSTC2024YCJH-BGZXM0022]; the Major Project of Science and Technology Research Program of Chongqing Education Commission of China [grant number KJZD-M202201102].